Income Or Output Y Consumption Expenditure         C Investment Expenditure       I Government Expenditure          G...

Income

Or

Output

Y

Consumption

Expenditure

        C

Investment

Expenditure

      I

Government

Expenditure

         G

Net export

Expenditure

       NX

$4,000

3,925

100

100

25

4,100

4,000

100

100

25

4,200

4,075

100

100

25

4,300

4,150

100

100

25

4,400

4,225

100

100

25

4,500

4,300

100

100

25

4,600

4,375

100

100

25

4,700

4,450

100

100

25

4,800

4,525

100

100

25

4,900

4,600

100

100

25

5,000

4,675

100

100

25

2. Calculate MPC, MPS and the Multiplier if consumption expenditure increases by $4,250 as a result of increase in income from $40,000 to $45,000.
Answer #1

MPC = change in consumption / Change in income

Here change or increase in consumption = 4250

Change or increase in income = 45000 - 40,000 = 5000

MPC = 4250 /5000 = 0.85

MPS = 1- MPC = 1- 0.85

MPS = 0.15

Multiplier = 1/ MPS = 1/0.15 = 6.67